OpenStudy (anonymous):

Which expression simplifies to 7h sqrt (2h)

OpenStudy (aripotta):

we need the choices brah

OpenStudy (anonymous):

ahh !!! forgots!

OpenStudy (anonymous):

%^e% is the answer.

OpenStudy (anonymous):

\[\sqrt{98h^3} \] \[\sqrt{14h^2}\] \[\sqrt{49h^3}\] \[\sqrt{49h^2}\]

OpenStudy (anonymous):

\[\sqrt{49h^2}\]

OpenStudy (aripotta):

that's not right. because that simplified would be 7h

OpenStudy (aripotta):

it's a

OpenStudy (anonymous):

jus square it, u ll find ur answer

OpenStudy (anonymous):

its sqrt(98h^3)?

geerky42 (geerky42):

7²·2 = 98

geerky42 (geerky42):

It is √98h³

OpenStudy (anonymous):

thanks!! :D

geerky42 (geerky42):

Remember, h = √h²

geerky42 (geerky42):

You're welcome.

OpenStudy (anonymous):

\[\sqrt{98h^3}\]This is the correct answer. This is because when you simplify it, this is what you get:\[\sqrt{98h^3}=\sqrt{49*h^2*2*h}=\sqrt{49h^2}\sqrt{2h}=7h \sqrt{2h}\]

OpenStudy (anonymous):

@vetdoc